ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 30F35 Aulaskari, Rauno, University of Joensuu Q_{p}-spaces on Riemann surfaces Recently defined function spaces Q_{p} are considered on hyperbolic Riemann surfaces.